3H decays with a rate constant of 0.0563/yr. What is the activity of a 0.699 g sample in curies?


1
Expert's answer
2021-04-25T06:15:34-0400

1 Ci = 2.79×10-6 × moles × t1/2

= 2.79×10-6 × "\frac{0.699}{3}" × "\frac{1n^2}{\lambda}"

= 2.79×10-6 × "\frac{0.699}{3}" × "\frac{0.693}{0.0563}"


= 8 × 10-6 yr